# Probability-Aware Parking Selection

### v1

@@ Abstract
Current parking navigation systems often underestimate total travel time by failing to account for the time spent searching for a parking space, which significantly affects user experience, mode choice, congestion, and emissions. To address this issue, this paper introduces the probability-aware parking selection problem, which aims to direct drivers to the best parking location rather than straight to their destination. An adaptable dynamic programming framework is proposed for decision-making based on probabilistic information about parking availability at the parking lot level. Closed-form analysis determines when it is optimal to target a specific parking lot or explore alternatives, as well as the expected time cost. Sensitivity analysis and three illustrative cases are examined, demonstrating the model’s ability to account for the dynamic nature of parking availability. Acknowledging the financial costs of permanent sensing infrastructure, the paper provides analytical and empirical assessments of errors incurred when leveraging stochastic observations to estimate parking availability. Experiments with real-world data from the US city of Seattle indicate this approach’s viability, with mean absolute error decreasing from 7% to below 2% as observation frequency grows. In data-based simulations, probability-aware strategies demonstrate time savings up to 66% relative to probability-unaware baselines, yet still take up to 123% longer than direct-to-destination estimates.
@@ I Introduction
Current navigation apps typically send drivers to their destination, occasionally providing a basic assessment (e.g., ‘easy’, ‘difficult’) of parking in the area [arora2019hard]. Given that parking spaces may not exist at the destination or may be unavailable, drivers often need to find parking elsewhere or cruise for parking at the current location until a spot becomes available. For example, consider the parking garage at San Francisco General Hospital, located in an area with limited on-street parking: when it is full, driving to and walking from the next-closest garage adds at least 30 minutes for patients and visitors, according to Google Maps routing estimates. Similar scenarios, familiar to many urban drivers and riders, highlight the frustrating difference between today’s ‘time to drive’ estimates that ignore parking difficulties and more useful ‘time to arrive’ estimates that account for true drive times and post-parking walk time.
@@
Beyond inconvenience, underestimating the true time to arrive via personal vehicles may cause two other problems. First, it may prevent mode shift that would occur if individuals had more accurate estimates [arora2019hard]. This is particularly true given that popular navigation apps (e.g., Google Maps, Apple Maps) include estimates for walking time from public transit to a final destination in their total public transit travel time estimates, but do not include walking time from available parking to a final destination in their total drive time estimates [Yafeng2024]. Second, cruising for parking is known to contribute to congestion and emissions in urban environments [shoup2006cruising].
@@
To address this issue, this paper introduces the probability-aware parking selection problem and proposes an adaptable dynamic programming framework to direct drivers to the best parking location based on location and parking lot-level probabilistic availability information, rather than directly to their destination. The dynamic programming framework offers a structured and adaptable theoretical foundation for closed-form analysis of optimal parking strategies under uncertainty, sensitivity, changing probabilities, as well as a mechanism for understanding trade-offs involved. Results leveraging real-world parking data from Seattle indicate stochastic observations of true lot-level parking probabilities could be a viable estimation approach. In the analyzed cases, observed probabilities’ mean absolute error begins below 7% and falls below 2% as observations become more frequent. Simulations with real-world data across different settings demonstrate the practical potential of probability-aware parking strategies. In all cases studied, these strategies outperform traditional methods, achieving time savings of up to 66% in the setting with the highest congestion. The work makes three contributions:
@@
  • An assessment of probability-aware parking strategies with stochastic observations, including a comparative case study using real-world parking data from Seattle.
The remainder of this work is organized as follows: Section II reviews the related literature. Section III introduces the dynamic programming problem and a corresponding solution, along with analysis for dynamic probability settings. Section IV details experiments that explore error from stochastic observations. Section V compares probability-aware and baseline search strategies. In Section VI the paper sets forth implications for future work.
@@ II Related Work
Probability-aware parking selection has key differences from the well-studied optimal stopping problem for parking [sakaguchi1982optimal, tamaki1982optimal, tamaki1988optimal], as well as other theoretical analyses of cruising [arnott2017cruising]. Primarily, the driver can decide where to go to seek parking, rather than passively waiting for availability to appear along a pre-defined route with strictly decreasing distance (usually) to the destination. Our analysis also considers probabilities at the level of entire parking lots rather than individual spaces, since the former is easier to estimate [xiao2023parking]. Still, our framework can accommodate spot-level analysis by treating each as its own individual lot. Even in optimal stopping variants that allow backtracking, vehicles cannot alter the order of their visits to various spaces, the driving geometry is 1-dimensional, and parking availability odds rarely vary [tamaki1988optimal, krapivsky2019simple].
Two well-executed closely related works are those described in [djuric2016parkassistant] and [hedderich2018optimization]. Given information about a driver’s parking preferences, the road network, traffic rules, and parking regulations, the algorithm in [djuric2016parkassistant] greedily calculates expected driver utility along possible routes to suggest a trajectory. [hedderich2018optimization] describes a routing approach based on the A* algorithm with a hand-designed cost function and focuses on the optimization of that function’s weight term, which is specific to a given road network and set of static parking probabilities. However, these works consider only street parking, do not use closed-form analysis to express or compare possibilities, and focus on cruising within a predefined tolerance boundary of the final destination instead of a priori selection of parking destinations separate from the final destination. Furthermore, neither formulates a dynamic programming framework for the problem, addresses how parking information may be obtained, nor considers dynamic parking probabilities.
Various mobile applications exist (e.g., SpotHero, SpotAngels, ParkCBR, Parking.com, Park Smarter) or have been proposed (e.g., ASPIRE [rizvi2018aspire]) to assist in parking, but generally these are simple reservation and/or payment systems for only those spots over which the app has control [teodorovic2006intelligent]. Other systems like iParker address pricing, but not probability-based lot selection [kotb2016iparker].
The survey in [xiao2023parking] as well as the works in [bock2019smart] and [shi2018parkcrowd] investigate parking availability predictions. The present paper builds upon this research to better understand the utility of such enhanced parking estimates for users. [wu2014agile] uses the possibility of time- and availability-aware parking selection to motivate the work, but focuses on empirical probability prediction at specific lots with in-built sensors in a vehicle’s immediate vicinity. Still, our notion of parking probability aligns with theirs, i.e., the likelihood that at least one parking space is available when a vehicle arrives.
@@ III-A Framework
The objective is to find the action sequence that maximizes the expected sum of rewards (equivalently, minimizes expected travel time). This framing excludes additional driver preferences such as pricing; future work could integrate these. The MDP can be seen as a stochastic shortest path problem in which the terminating goal states $G\subset\mathcal{S}$ are the parked states; this termination condition obviates the need for a discount factor $\gamma$ in the framing.
III-B Optimal strategy and time cost in the probabilistic information setting
This work focuses on the case in which probabilistic parking availability information is accessible for each lot since it most closely relates to today’s app-based driving decisions. Given pre-existing awareness of parking difficulty and app-generated data, it is uncommon to truly have no prior information about parking availability odds [arora2019hard]. At the same time, real-time parking occupancy sensors are expensive and thus relatively rare [djuric2016parkassistant].
@@ III-B Optimal strategy and time cost in the probabilistic information setting
Note that the decision to park at lot $i$ can be conceptualized as a coin flip that succeeds (allows parking) with probability $p_{i}$. $t_{\text{wait}}$ can be considered the time one must wait at a lot to again ‘flip the coin’ if the previous parking attempt was unsuccessful. Adopting a dynamic programming approach, one can thus explicitly write the cost for any ‘patient’ strategy, i.e., the strategy in which a vehicle drives to the $i$th lot and waits there until it finds parking (possible since $p_{i}\in(0,1]$). The expected cumulative return $\mathbf{E}[R_{i,\text{patient}}]$ of this strategy is written as
@@
While this may seem a strong assumption, drive times to parking lots near a destination of interest are often quite comparable, especially relative to drive time variance due to other factors, e.g., traffic congestion, weather [peer2012prediction]. As an indication of this, empirical distributions of drive times on urban roads have been shown to exhibit standard deviations one-third of their means [fosgerau2012valuing]. Since we are concerned with a priori destination selection, this variance enhances the extent to which drive times to two nearby locations may appear functionally similar, especially in comparison to the other time costs involved, thus expanding the proposition’s scope.
@@ III-D Dynamic probabilities & three illustrative cases
Given the above sensitivity analysis, we now consider dynamic probabilities. In many settings, we do not expect walk time estimates between a parking lot and a destination of interest to vary considerably, and estimating driving time is a well-known feature of many of today’s navigation apps as well as the subject of a number of previous works [bertsimas2019travel, uchida2014estimating, wang2018will]. Conversely, the probability of parking at a specified lot may vary considerably throughout the day due to demand, duration of individual stays at a location, etc.
@@
  • Third-order: Another vehicle does not attempt to park in the ego vehicle’s lot, but nevertheless affects the ego vehicle’s parking probability via impacts on an intervening parking lot.
@@ III-E Errors from stochastic probability observations
If the underlying lot-level probabilities are dynamic, stochastic measurements will incur error as their recency fades. This can be an issue in light of the sensitivity analysis above; i.e., if the observed probabilities deviate too much from the true probabilities, vehicles might route to suboptimal locations. Thus, we provide analysis and experiments below to investigate errors incurred when estimating dynamic probabilities via Poisson process-generated observations. Poisson processes are a well-known stochastic modeling tool for random arrival processes, including for parking analysis [newell1982, Jorge2012, kutoyants2023, xiao2018likely, ogulenko2022nature].
The analysis below applies to any pointwise-independent stochastic observation method and could easily be adapted to a variety of schemes. [xiao2023parking] provides a comprehensive survey on parking prediction, particularly via data-driven methods for probability prediction. Near-term prediction is also discussed in [xiao2018likely]. For concreteness – and given the costs and effort associated with permanent infrastructure, continuous remote sensing, etc. – the current work considers that these Poisson observations are generated by connected users, where such users represent $r$ fraction of the total population. This is inspired by the approach in [shi2018parkcrowd]. For example, users of a given routing app or connected vehicles (autonomous or not) could provide insight into the availability at a lot to which they arrive or pass. [bock2019smart] has illustrated the feasibility of a similar method, although that work leverages taxis and considers only on-street parking. Note, however, that it would be straightforward to adapt the analysis to other Poisson stochastic observation methods, including schemes with remote sensing or paid staff, by adjusting the rate parameter.
@@ V-A1 Sites
To validate the results of the proposed estimation approach across different times of day and urban parking settings, we introduce data provided by Seattle’s Department of Transportation (SDOT) [seattle_paid_parking_occupancy, seattle_paid_parking_transactions] for two separate areas:
@@
They are selected to reflect heterogeneity: Area A.1 is a dense urban setting in the heart of downtown Seattle with high traffic, while Area B.1 is located five miles northwest and represents an area with lighter traffic, less dense zoning, and sparser SDOT parking options.
@@ V-A2 Scenario
We employ payment transactions collected on Thursday, January 30th, 2025 as evidence of arrivals and take inverse occupancy rates as a proxy for parking probability. The data is available for Area A.1 from 8am to 8pm and for Area B.1 from 8am to 6pm. Occupancy is computed from paid parking occupancy and the total parking space count. To simulate connected users, we randomly sample the transaction data and assign the appropriate proportion as connected user arrivals.
@@ V-A3 Results
We evaluate stochastic probability observation via connected users by computing errors as in Section IV. Results are shown in Figure 6 with the same conventions as in Figures 4 and 5. Given the connected user sampling involved, for the error box plots we again evaluated 100 random seeds at each adoption rate. In all cases the mean and median MAE is below 7% and declines with increased adoption to below 2%. Interestingly, while Figure 6a shows the relative discrepancy between true and observed probabilities in Area B.1 to be greater than that at Area A.1 due to the lower effective arrival rate of connected users, Figure 6b – by the same token – indicates that the lower volatility in Area B.1 parking probability also produces MAEs that are more consistently low.
@@
TABLE I: Comparison of Probability-Aware Selection Policies with Patient and Impatient Baselines Across Temporal and Spatial Settings.
@@
BL-Impat. 13.8 ± 5.2 5.1% 14.0 ± 5.5 -0.1%
PA 1-Step 13.4 ± 4.5 8.1% 3.2% 1.1% 11.6 ± 2.4 17.1% 17.3% 5.5%
PA 2-Step 13.5 ± 3.6 7.3% 2.4% 1.3% 13.4 ± 3.4 4.4% 4.6% -5.1%
PA 3-Step 13.8 ± 5.5 5.1% 0.0% -8.3% 13.6 ± 5.4 3.0% 3.1% -9.7%
For brevity, ‘PA’ indicates probability-aware policies and ‘BL’ indicates baselines. Up arrows indicate that higher values are preferable for the metric, while down arrows indicate the opposite. Times are reported in minutes. Boldface denotes the top performance within each adoption rate category. See the main text for further details.

### v2

@@ Abstract
Current navigation systems conflate time-to-drive with the true time-to-arrive by ignoring parking search duration and the final walking leg. Such underestimation can significantly affect user experience, mode choice, congestion, and emissions. To address this issue, this paper introduces the probability-aware parking selection problem, which aims to direct drivers to the best parking location rather than straight to their destination. An adaptable dynamic programming framework is proposed that leverages probabilistic, lot-level availability to minimize the expected time-to-arrive. Closed-form analysis determines when it is optimal to target a specific parking lot or explore alternatives, as well as the expected time cost. Sensitivity analysis and three illustrative cases are examined, demonstrating the model’s ability to account for the dynamic nature of parking availability. Given the high cost of permanent sensing infrastructure, we assess the error rates of using stochastic observations to estimate availability. Experiments with real-world data from the US city of Seattle indicate this approach’s viability, with mean absolute error decreasing from 7% to below 2% as observation frequency increases. In data-based simulations, probability-aware strategies demonstrate time savings up to 66% relative to probability-unaware baselines, yet still take up to 123% longer than time-to-drive estimates.
@@ I Introduction
Current navigation apps typically send drivers to their destination, occasionally providing a basic assessment (e.g., ‘easy’, ‘difficult’) of parking in the area [2]. Given that parking spaces may not exist at the destination or may be unavailable, drivers often need to find parking elsewhere or cruise for parking at the current location until a spot becomes available. For example, consider the parking garage at San Francisco General Hospital, located in an area with limited on-street parking: when it is full, driving to and walking from the next-closest garage adds at least 30 minutes for patients and visitors, according to Google Maps routing estimates. Similar scenarios, familiar to many urban drivers and riders, highlight the frustrating difference between today’s time-to-drive estimates that ignore parking difficulties and more useful time-to-arrive estimates that account for the parking search and post-parking walk time. See Figure 1.
Figure 1: Accounting for the search for parking and the post-parking walking leg can increase the accuracy of personal vehicle travel time estimates, save time, and improve cross-mode comparability.
@@
Beyond inconvenience, underestimating the true time to arrive via personal vehicles may cause two other problems. First, it may prevent mode shift that would occur if individuals had more accurate estimates [2]. This is particularly true given that popular navigation apps (e.g., Google Maps, Apple Maps) include estimates for walking time from public transit to a final destination in their total public transit travel time estimates, but do not include walking time from available parking to a final destination in their total drive time estimates [29]. Second, cruising for parking is known to contribute to congestion and emissions in urban environments [20].
@@
To address this issue, this paper introduces the probability-aware parking selection problem and proposes an adaptable dynamic programming framework to direct drivers to the best parking location based on location and parking lot-level probabilistic availability information, rather than directly to their destination. The dynamic programming framework offers a structured and adaptable theoretical foundation for closed-form analysis of optimal parking strategies under uncertainty, sensitivity, changing probabilities, as well as a mechanism for understanding trade-offs involved. Results leveraging real-world parking data from Seattle indicate stochastic observations of true lot-level parking probabilities could be a viable estimation approach. In the analyzed cases, observed probabilities’ mean absolute error begins below 7% and falls below 2% as observations become more frequent. Simulations with real-world data across different settings demonstrate the practical potential of probability-aware parking strategies. In all cases studied, these strategies outperform traditional methods, achieving time savings of up to 66% in the setting with the highest congestion. The work makes four contributions:
  • It establishes time-to-arrive as a unified metric to enable fair comparison across transport modes, correcting the bias inherent in omitting search and walk times.
@@
  • A demonstration of the time savings from probability-aware parking strategies with stochastic observations. This includes a comparative case study using real-world parking data from Seattle that illustrates the utility of time-to-arrive relative to time-to-drive.
This work is organized as follows: Section II reviews the related literature. Section III introduces the dynamic programming problem and a corresponding solution, along with analysis for dynamic probability settings. Section IV details experiments that explore error from stochastic observations. Section V presents the Seattle case study. In Section VI the paper sets forth implications for future work.
@@ II Related Work
Probability-aware parking selection has key differences from the well-studied optimal stopping problem for parking [18, 21, 22], as well as other theoretical analyses of cruising [1]. Primarily, the driver can decide where to go to seek parking, rather than passively waiting for availability to appear along a pre-defined route with strictly decreasing distance (usually) to the destination. Our analysis also considers probabilities at the level of entire parking lots rather than individual spaces, since the former is easier to estimate [28]. Still, our framework can accommodate spot-level analysis by treating each as its own individual lot. Even in optimal stopping variants that allow backtracking, vehicles cannot alter the order of their visits to various spaces, the driving geometry is 1-dimensional, and parking availability odds rarely vary [22, 9].
Two well-executed closely related works are those described in [5] and [7]. Given information about a driver’s parking preferences, the road network, traffic rules, and parking regulations, the algorithm in [5] greedily calculates expected driver utility along possible routes to suggest a trajectory. [7] describes a routing approach based on the A* algorithm with a hand-designed cost function and focuses on the optimization of that function’s weight term, which is specific to a given road network and set of static parking probabilities. However, these works consider only street parking, do not use closed-form analysis to express or compare possibilities, and focus on cruising within a predefined tolerance boundary of the final destination instead of a priori selection of parking destinations separate from the final destination. Furthermore, neither formulates a dynamic programming framework for the problem, addresses how parking information may be obtained, nor considers dynamic parking probabilities.
Various mobile applications exist (e.g., SpotHero, SpotAngels, ParkCBR, Parking.com, Park Smarter) or have been proposed (e.g., ASPIRE [17]) for parking, but generally these are simple reservation and/or payment systems only for spots over which the app has control [23]. Other systems like iParker address pricing, but not probability-based selection [8].
The survey in [28] as well as the works in [4] and [19] investigate parking availability predictions. The present paper builds upon this research to better understand the utility of such enhanced parking estimates for users. [26] uses the possibility of time- and availability-aware parking selection to motivate the work, but focuses on empirical probability prediction at specific lots with in-built sensors in a vehicle’s immediate vicinity. Still, our notion of parking probability aligns with theirs, i.e., the likelihood that at least one parking space is available when a vehicle arrives.
@@ III-A Framework
[…]
This distinction is critical for fair comparison across transportation modes. Current navigation platforms present time-to-drive estimates alongside time-to-arrive mass transit estimates that include wait and access/egress walk times. By omitting search and walking components, standard driving estimates present a structurally optimistic bias. Users, however, may compare app-displayed values across modes as equivalent estimates for total travel time. By explicitly modeling and optimizing for time-to-arrive, we correct this asymmetry, enabling valid cross-mode comparisons.
The objective is therefore to find the action sequence that minimizes expected time-to-arrive (equivalently, maximizes the expected sum of rewards). This framework excludes additional driver preferences, such as pricing; future work could integrate these. The MDP can be seen as a stochastic shortest path problem in which the terminating goal states $G\subset\mathcal{S}$ are the parked states; this termination condition obviates the need for a discount factor $\gamma$ in the framing.
III-B Optimal strategy and cost with probabilistic information
This work focuses on the case in which probabilistic parking availability information is accessible for each lot since it most closely relates to today’s app-based driving decisions. Given pre-existing awareness of parking difficulty and app-generated data, it is uncommon to truly have no prior information about parking availability odds [2]. At the same time, real-time parking occupancy sensors are expensive and thus rare [5].
@@ III-B Optimal strategy and cost with probabilistic information
Note that the decision to park at lot $i$ can be conceptualized as a coin flip that succeeds (allows parking) with probability $p_{i}$. $t_{\text{wait}}$ can be considered the time one must wait at a lot to again ‘flip the coin’ if the previous attempt was unsuccessful. Adopting a dynamic programming approach, one can thus explicitly write the cost for any ‘patient’ strategy, i.e., the strategy in which a vehicle drives to the $i$th lot and waits there until it finds parking (possible since $p_{i}\in(0,1]$). The expected cumulative return $\mathbf{E}[R_{i,\text{patient}}]$ of this strategy is written as
@@
While this may seem a strong assumption, drive times to parking lots near a destination of interest are often quite comparable, especially relative to drive time variance due to other factors, e.g., traffic congestion, weather [16]. As an indication of this, empirical distributions of drive times on urban roads have been shown to exhibit standard deviations one-third of their means [6]. Since we are concerned with a priori destination selection, this variance enhances the extent to which drive times to two nearby locations may appear functionally similar, especially in comparison to the other time costs involved, thus expanding the proposition’s scope.
@@ III-D Dynamic probabilities & three illustrative cases
Given the above sensitivity analysis, we now consider dynamic probabilities. In many settings, we do not expect walk time estimates between a parking lot and a destination of interest to vary considerably, and estimating driving time is a well-known feature of many of today’s navigation apps as well as the subject of a number of previous works [3, 24, 25]. Conversely, the probability of parking at a specified lot may vary considerably throughout the day due to demand, duration of individual stays at a location, etc.
@@
  • Third-order: Another vehicle does not attempt to park in the ego vehicle’s lot, but affects the ego vehicle’s parking probability via impacts on an intervening parking lot.
@@ III-E Errors from stochastic probability observations
If the underlying lot-level probabilities are dynamic, stochastic measurements will incur error as their recency fades. This can be an issue in light of the sensitivity analysis above; i.e., if the observed probabilities deviate too much from the true probabilities, vehicles might route to suboptimal locations. Thus, we provide analysis and experiments below to investigate errors incurred when estimating dynamic probabilities via Poisson process-generated observations. Poisson processes are a well-known stochastic modeling tool for random arrival processes, including for parking analysis [12, 11, 10, 27, 15].
The analysis below applies to any pointwise-independent stochastic observation method and could easily be adapted to a variety of schemes. [28] provides a comprehensive survey on parking prediction, particularly via data-driven methods for probability prediction. Near-term prediction is also discussed in [27]. For concreteness – and given the costs and effort associated with permanent infrastructure, continuous remote sensing, etc. – the current work considers that these Poisson observations are generated by connected users, where such users represent $r$ fraction of the total population. This is inspired by the approach in [19]. For example, users of a given routing app or connected vehicles (autonomous or not) could provide insight into the availability at a lot to which they arrive or pass. [4] has illustrated the feasibility of a similar method, although that work leverages taxis and considers only on-street parking. Note, however, that it would be straightforward to adapt the analysis to other Poisson stochastic observation methods, including schemes with remote sensing or paid staff, by adjusting the rate parameter.
@@ V Case study: Seattle parking
Figure 7: Locations of the destinations used in the Seattle case study. Base map sources: MapTiler, OpenStreetMap contributors.
@@ V-A1 Sites
To validate the results of the proposed estimation approach across different times of day and urban parking settings, we introduce data provided by Seattle’s Department of Transportation (SDOT) [13, 14] for two separate areas:
@@
They are selected to reflect heterogeneity: Area A.1 is a dense urban setting in the heart of downtown Seattle with high traffic, while Area B.1 is located five miles northwest and represents an area with lighter traffic, less dense zoning, and sparser SDOT parking options. See Figure 7.
@@ V-A2 Scenario
We use payment transactions collected on Thurs., Jan. 30th, 2025 as evidence of arrivals and take inverse occupancy rates as a proxy for parking probability. The data is available for Area A.1 from 8am to 8pm and for Area B.1 from 8am to 6pm. Occupancy is computed from paid parking occupancy and the total parking space count. To simulate connected users, we randomly sample the transaction data and assign the appropriate proportion as connected user arrivals.
@@ V-A3 Results
We evaluate stochastic probability observation via connected users by computing errors as in Section IV. Results are shown in Figure 8 with the same conventions as in Figures 5 and 6. Given the connected user sampling involved, for the error box plots we again evaluated 100 random seeds at each adoption rate. In all cases the mean and median MAE is below 7% and declines with increased adoption to below 2%. Interestingly, while Figure 8a shows the relative discrepancy between true and observed probabilities in Area B.1 to be greater than that at Area A.1 due to the lower effective arrival rate of connected users, Figure 8b indicates that – by the same token – the lower resulting volatility in Area B.1’s parking probability produces MAEs that are more consistently low.
@@
TABLE I: Comparison of Probability-Aware Selection Policies with Patient and Impatient Baselines Across Temporal and Spatial Settings
@@
[…]
For brevity, ‘PA’ indicates probability-aware policies and ‘BL’ indicates baselines. Up arrows indicate that higher values are preferable for the metric, while down arrows indicate the opposite. Times are reported in minutes. Boldface denotes the top performance within each adoption rate category. See the main text for further details. Darker shading indicates larger performance differences.